Accelerator Update for December 8, 2008 - December 10, 2008
Monday Dec 8
The day shift began with Operations monitoring the Tevatron (TeV) store 6628 (114.22E30) and the Recycler stash (291.4E10), with the Antiproton Source stacking (33.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the MiniBooNE experiment taking beam. NuMI was down due to a horn problem.
Operations terminated store 6628 at 10:59 AM.
Operations put the TeV into shot setup at 11:09 AM.
At 11:26 AM, Operations reported that the I- Source was sparking. Linac experts investigated.
Operations established store 6629 at 12:52 PM with an initial luminosity of 333.72E30.
A Pbar expert began an AP-10 study at 2:44 PM. This study disrupted stacking.
A MI expert began a multi- batch study at 3:28 PM.
Operations reported at 4:01 PM that the Recycler had stashed antiprotons four times during the day shift.
The MI study ended at 4:19 PM.
The Pbar expert completed his AP-10 study at 5:10 PM.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Tuesday Dec 9
The midnight shift began with Operations monitoring store 6629 (97.99E30) and the stash (200.8E10), with Pbar stacking (29.7E10), and with MiniBooNE taking beam.
Operations terminated store 6629 around 7:17 AM.
Operations put the TeV into shot setup at 7:53 AM.
Operations reported at 7:59 AM that the Recycler had stashed eight times during the midnight shift.
Operations reported at 8:41 AM that a Linac expert adjusted the timing for some of the Linac quadrupoles due to a noise problem.
At 9:09 AM, Operations reported an ODH alarm at MI-31. Operators instructed all the people in the service building to leave. An expert investigated and believed the alarm was due to a bad oxygen sensor in the tunnel. He would need an access to replace the bad sensor.
Operations established store 6630 at 9:23 AM with an initial luminosity of 325.37E30.
A MI expert began a study at 9:53 AM.
The MI expert completed her study at 11:06 AM.
At 11:47 AM, Operations turned off Linac for work on the quadrupole power supplies. Operations put Pbar into standby mode.
The Linac work ended at 12:32 PM. Pbar resumed stacking.
TeV experts began a one-hour study around 2:06 PM.
Operations reported at 3:58 PM that the Recycler had stashed five times during the day shift.
At 4:25 PM, the P1 transfer line beam permit tripped off when the current from a power supply (I:V701A) drifted. It reset.
At 4:29 PM, Operations reported more problems with V701A. Operators turned off the supply and called in experts. Transfers of antiprotons to the Recycler were put on hold.
At 6:15 PM, the P1 transfer line tripped off for the first of six times in 52 minutes. It reset.
Power supply experts reported at 6:25 PM that when they put a scope on the supply to troubleshoot the problem – the problem went away. They reported that this probably pointed to a problem with an internal 480-volt breaker. They staged a spare breaker.
At 6:34 PM, all the Pbar power supplies at AP2 and AP3 tripped off due to a bad LCW flow indication. Operations called in an expert.
At 7:29 PM, LCW experts reported that the Pbar LCW pump for the F27 closed loop water skid had seized bearings. They began work on replacing the pump.
At 7:44 PM, V701A began causing problems again. A power supply expert investigated.
A TeV expert began a Tevatron Electron Lens measurement at 8:15 PM.
At 9:57 PM, a power supply expert reported that he was not sure if the problem was with V701A or B or even what the problem was. He installed a transient recorder.
At 9:59 PM, LCW experts reported that they had replaced the water skid pump, but found a wiring problem.
The Duty Electrician reported that he had completed Pbar pump wiring. Pbar resumed stacking at 11:35 PM.
A power supply expert reported at 11:39 PM that he had narrowed the problem with the V701 supply to two boards. He left both boards and instructions for the operators on how to replace them when the power supply failed again.
Operations reported at midnight that the Recycler had stashed once during the evening shift.
Wednesday Dec 10
The midnight shift began with Operations monitoring store 6630 (77.28E30) and the stash (212.5E10), with Pbar stacking (21.9E10), and with MiniBooNE taking beam.
At 12:42 AM, Operations reported two vacuum bursts in the MI-60 area that caused all the RF stations to trip off. Everything reset.
At 12:57 AM, Operations reported that there was still a MI vacuum problem in the MI-60 area.
At 1:50 AM, Operations reported sequencer problems with the Pbar transfers of antiprotons to the Recycler. Operators investigated and discovered a stacktail medium level amplifier had tripped off. It reset.
Operations terminated store 6630 at 3:09 AM.
Operations put the TeV into shot setup at 3:22 AM.
At 4:07 AM, the Linac Buncher tripped off. It reset.
Operations established store 6631 at 5:10 AM with an initial luminosity of 275.16E30.
Operations reported at 6:17 AM that the vacuum had recovered in the MI-60 area.
Operations reported at 8:01 AM that the Recycler had stashed 5 times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator will allow TeV experts to conduct a crystal collimator study this afternoon.
Comments and Suggestions
|last modified 01/05/2009 email Fermilab|
|Security, Privacy, Legal|